OpenStudy (anonymous):

x/5-8=12 2(3x+1) ------- = -8 -4

OpenStudy (anonymous):

x/5-8=12 x/5=20 x=100 2(3x+1) -------= -8 -4 2(3x+1)=32 3x+1=16 3x=15 x=5